# devoted to memory of Alexei Alexandrovich Starobinsky

### v15

@@ V.2 Entropy of expanding, contracting and static de Sitter
This metric has singularity at the cosmological horizon, which requires the proper attention. In one approach this static configuration can be considered as the intermediate state between the two opposite states – the expanding de Sitter state with $H>0$ and the contracting de Sitter state with $H<0$ (see the same consideration of the static hole as the intermediate state between black and white holes in Ref. Volovik2022 ). In the contracting de Sitter vacuum with $H<0$ the local temperature is negative, $T=H/\pi<0$, and the local entropy is also negative, $s_{\rm contracting}=3H/4G=-3/(4Gr_{h})<0$. The total entropy in the Hubble volume of contracting de Sitter is
@@
where $A$ is the area of the cosmological horizon. That is why the entropy of the static de Sitter as the intermediate state is zero, and correspondingly the temperature must be infinite, see also Ref.Milekhin2024 .
V.3 Gravastar – black hole with de Sitter core
Let as consider the black hole, which is obtained by the deformation of the gravastar.Volovik2023a The gravastar is the object, which contains the de Sitter state inside the black hole horizon.Chapline2003 ; Mazur2023 ; Mottola2023 We consider the gravastar in which the black hole horizon coincides with the de Sitter horizon, $r_{\rm bh}=r_{h}$ (the metric in the state with the critical value of the mass parameter $m=2MG|H|=1$ at which two horizons merge is illustrated in Fig. 1 of Ref. Dymnikova2002 ). Such gravastar is stationary, that is why it has no Hawking radiation and its entropy is zero. In the Painlevé-Gullstrand form the metric of such gravastar is given by Eq.(1) with the following shift velocity:
@@ V.3 Gravastar – black hole with de Sitter core
The shift velocity $v(r)$ must be continuous across the horizon, i.e. there is no jump in the shift velocity, while the gradient of the shift velocity $dv/dr$ experiences jump at the horizon. The shift velocity $v(r)$ is negative both outside and inside the horizon. Since the shift velocity is negative in the de Sitter core, this means that the de Sitter spacetime in this gravastar is contracting, $v(r)=Hr=-r/r_{h}<0$, i.e. the Hubble parameter is negative, $H=-1/r_{h}<0$.
V.4 Entropy of black hole from negative entropy of contracting de Sitter
@@
V.5 White hole and anti-gravastar
@@
It is important that the singularity in the metric of gravastar and in the metric of anti-gravastar can be smoothly removed by deformation. That is why these objects do not depend on the choice of the coordinate systems, and thus they are equivalent to the fully static black hole with the fully static de Sitter core:
@@
V.6 Heat exchange between black holes in the multi-metric ensemble
Let us first consider the two-metric gravity with $K_{1}=K_{2}$. In this case, the heat exchange between the black holes with masses $M_{1}$ and $M_{2}$ at fixed $M=M_{1}+M_{2}$ leads to formation of the black holes with masses $M_{1}=M_{2}=M/2$ in each of the two sub-Universes. For sub-Universes with $K_{1}\neq K_{2}$ one obtains $M_{1}=MK_{1}/(K_{1}+K_{2})$ and $M_{2}=MK_{2}/(K_{1}+K_{2})$.
For the general ensemble of the multi-metric gravities, the heat exchange between the substates gives rise to the black hole with mass $M$ and gravitational coupling $K$:
$M=\sum_{n}M_{n}\,\,,\,\,K=\sum_{n}K_{n}\,\,,\,\,M_{n}=M\,\frac{K_{n}}{K}\,.$ (48)
The Hawking temperature of the black hole is the common temperature in all the substates:

### v16

@@ VII.2 Entropy of expanding, contracting and static de Sitter
This metric has singularity at the cosmological horizon, which requires the proper attention. In one approach this static configuration can be considered as the intermediate symmetric state between the two states with broken time reversal symmetry – the expanding and contracting states. The same consideration applies to the static hole, that can be viewed as the intermediate state between the black and white holes.Volovik2022 .
The contracting de Sitter vacuum has negative Hubble parameter, $H<0$. That is why its local temperature is negative, $T=H/\pi<0$, and the local entropy is also negative, $s_{\rm contracting}=3H/4G=-3/(4Gr_{H})<0$. The total entropy in the Hubble volume of contracting de Sitter is
@@
where $A$ is the area of the cosmological horizon. That is why the entropy of the static de Sitter in Eq.(76) as the intermediate state between the states with positive and negative entropies is zero. Correspondingly the temperature of this state must be infinite, which is consistent with the singularity at the horizon, see also Ref.Milekhin2024 .
VII.3 Gravastar – black hole with de Sitter core
The connection between the black hole and de Sitter appears, when we consider the black hole obtained by the deformation of the gravastar.Volovik2023a The gravastar is an object, which contains the de Sitter vacuum inside the black hole horizon.Chapline2003 ; Mazur2023 ; Mottola2023 We consider the gravastar in which the black hole horizon coincides with the de Sitter horizon, $r_{\rm bh}=r_{H}$ (the metric in the state with the critical value of the mass parameter $m=2MG|H|=1$ at which two horizons merge is illustrated in Fig. 1 of Ref. Dymnikova2002 ). Since in such gravastar the two horizons cancel each other, there is no Hawking radiation and the entropy of the gravastar is zero. In the Painlevé-Gullstrand form the metric of such gravastar is given by Eq.(1) with the following shift velocity:
@@ VII.3 Gravastar – black hole with de Sitter core
The shift velocity $v(r)$ is continuous across the surface $r=r_{H}$, while the gradient of the shift velocity $dv/dr$ experiences jump at this surface. It is important that the shift velocity $v(r)$ is negative everywhere. Since it is negative in the de Sitter core, this means that the de Sitter spacetime in the core of this gravastar is contracting, $v(r)=Hr=-r/r_{H}<0$, i.e. the Hubble parameter is negative, $H=-1/r_{H}<0$.
VII.4 Entropy of black hole from negative entropy of contracting de Sitter
@@
VII.5 White hole and anti-gravastar
@@
It is important that the coordinate singularities in the metric of the gravastar and in the metric of the anti-gravastar can be smoothly removed by small deformations. That is why these objects do not depend on the choice of the coordinate systems, and thus they are equivalent to the fully static black hole with the fully static de Sitter core:
@@
This again supports the zero entropy of these gravastars, and also the zero entropy of the fully static de Sitter with metric in Eq.(76).
VII.6 Gibbs-Duhem and black hole thermodynamics
Let us show, that the modified Gibbs-Duhem relation in Eq.(46) is applicable also to the thermodynamics of black holes. As distinct from the de Sitter state, the black hole is the compact object, and its thermodynamics is operating with the global parameters, such as mass $M$, entropy of horizon, total electric charge $Q$ and total angular momentum $J$. This global thermodynamics can be described by the integral form of the Gibbs-Duhem relation in Eq.(46). The curvature here comes from the central singularity:Balasin1993
${\cal R}=8\pi MG\,\delta({\bf r})\,.$ (87)
Since the energy density here is $\epsilon=M\delta({\bf r})$, the integration of the right-hand-side of Eq.(46) over space gives the following relation for the Schwarzschild black hole:
$T_{\rm BH}S_{\rm BH}=M-\int d^{3}r\sqrt{-g}K{\cal R}=\frac{M}{2}\,.$ (88)
This agrees with global thermodynamics of the black hole with the Hawking temperature $T_{\rm BH}=\frac{1}{8\pi MG}$ and the Bekenstein-Hawking entropy $S_{\rm BH}=A/4G$. The Eq.(88) is valid also for the white hole, where temperature and entropy are opposite to that of the black hole with the same mass,Volovik2022 $T_{\rm WH}(M)=-T_{\rm BH}(M)$ and $S_{\rm WH}(M)=-S_{\rm BH}(M)$.
In principle, the central singularity in the black hole may spontaneously loose the spherical symmetry, forming for example a kind of rigid top. This in turn may influence the shape of the event horizon. This is similar to the deformations of the cosmological horizon by masses concentrated on vertices of Platonic solids deep within the Hubble volume.Fischler2024
VII.7 Entropy of the Schwarzschild-de Sitter cosmological horizon
Let us consider the possible application of the modified Gibbs-Duhem relation to the Schwarzschild-de Sitter (SdS) black hole. We discuss the simple case of the Nariai limit, when the black hole horizon approaches the cosmological horizon, $r_{b}\rightarrow r_{0}-0$ and $r_{c}\rightarrow r_{0}+0$, where
$r_{0}=(GM/H^{2})^{1/3}=\frac{1}{\sqrt{3}H}\,.$ (89)
In this limit the temperatures of the horizons approach the Bousso-Hawking value:BoussoHawking1996
$T_{b}=T_{c}=\frac{\sqrt{3}}{2\pi}H=\frac{1}{6\pi GM}\,.$ (90)
The entropy of the cosmological horizon $S_{\rm c}$ can be obtained by integration of the right-hand-side of Eq.(46) over the vollume inside the cosmological horizon with radius $r_{0}$:
$T_{\rm c}S_{\rm c}=M-\int_{r<r_{0}}d^{3}r\sqrt{-g}K{\cal R}\,.$ (91)
Since we discuss here the Nariai limit, one obtains:
$T_{\rm c}S_{\rm c}=M-\int_{r<r_{0}}d^{3}r\sqrt{-g}K{\cal R}=M-\frac{M}{2}+\frac{3H^{2}}{4\pi}\,\frac{4\pi r_{0}^{3}}{3}=\frac{3}{2}M.$ (92)
From this equation one obtains the entropy of the cosmological horizon:
$S_{\rm c}=\frac{3M}{2T_{\rm c}}=\frac{\pi r_{0}^{2}}{G}=\frac{A}{4G}\,.$ (93)
This again agrees with the Gibbons-Hawking entropy attributed to the horizon, although this holographic connection is valid only in the Narai limit.
VII.8 Heat exchange between black holes in the multi-metric ensemble
The thermodynamic character of the black hole objects can be tested using the multi metric ensemble in Sec. IV. Let us consider ensemble of $n$ sub-universes in the same coordinate spacetime, $S=-\int d^{4}x\sum_{n}{\cal L}_{n}$, but with with different metrics $g_{\mu\nu(n)}$, different gravitational couplings $K_{n}$, and the corresponding individual black holes with masses $M_{n}$. If one introduces the energy exchange between these sub-universes, then the masses $M_{n}$ will be varied at the fixed total mass $M=\sum_{n}M_{n}$ of the whole Universe. In the thermal equilibrium, which corresponds to the maximum of the total entropy, the individual metrics approach the common metric, and the whole system corresponds to the Universe with the gravitational coupling $K=\sum_{n}K_{n}$ and with the black hole of mass $M$. The equilibrium masses $M_{n}$ of the individual substates approach the values:
$M_{n}=M\,\frac{K_{n}}{K}\,.$ (94)
In this equilibrium states, the Hawking temperature of the black hole with mass $M$ in the whole ensemble coincides with Hawking temperatures of the individual black holes in each gravitational sub-states: